# Riemannian Optimization for Distance Geometry: A Study of Convergence, Robustness, and Incoherence

### v1

@@ 2.2 Dual Basis Approach to EDG
It can also easily be verified that the dimension of the linear space $\mathbb{S}$ is $L=n(n-1)/2$. Using this basis, we can realize each entry of the squared distance matrix as the trace inner product of the Gram matrix with the basis. Formally, $D_{ij}=\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle$ for $\bm{\alpha}=(i,j)$. Further, we can introduce the dual basis to $\{\bm{w}_{\bm{\alpha}}\}$, denoted as $\{\bm{v}_{\bm{\alpha}}\}$, and represent any centered Gram matrix $\bm{X}$ using the following expansion:
@@
The advantage of the dual basis representation is that it allows us to recast the EDG problem as a low-rank matrix recovery problem where we observe a subset of the expansion coefficients. In [28], this dual basis formulation has been used to provide theoretical guarantees for the convex program given in (5).
@@
The bi-orthogonality relationship of the dual basis gives that $\mathcal{R}_{\Omega}^{2}=\mathcal{R}_{\Omega}$ if $\Omega$ does not have repeated indices, and that
@@
3 Geometric Interpretation of EDG Incoherence
In pathological cases, the ground truth matrix $\bm{X}$ may exhibit a sparse representation in the basis $\{\bm{w}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$, which could lead to challenges in its recovery from sampled measurements. While the concept of incoherence is well-established in the standard matrix completion literature, the condition specific to the EDG problem slightly differs in structure and admits a natural geometric interpretation. This section is devoted to a detailed examination of this geometric perspective. We will state more formally the incoherence assumptions in Section 5, but we will first introduce one of the conditions below. We say that a $r$ Gram matrix $\bm{X}\in\mathbb{R}^{n\times n}$ is $\nu$-incoherent with respect to $\{\bm{w}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$ if the following statement holds:
@@ 3 Geometric Interpretation of EDG Incoherence
We remark that the above is inspired by the standard incoherence condition, which states that
@@
The incoherence condition introduced in (11) can be interpreted in terms of the underlying point cloud data. For the specific case of the EDG problem, (11) can be expanded as follows for $\bm{\alpha}=(i,j)$ with $i<j$:
@@
We consider $\sum_{(i,j):i<j}\left(\bm{u}_{i}-\bm{u}_{j}\right)^{\top}\left(\bm{u}_{i}-\bm{u}_{j}\right)$. Note that $\sum_{i}(\bm{u}_{i})^{\top}\bm{u}_{i}=\mathrm{Trace}(\bm{U}\bm{U}^{\top})=\mathrm{Trace}(\bm{U}^{\top}\bm{U})=r$. Since we assume centered configurations, $\bm{U}^{T}\bm{1}=\bm{0}$. It then follows that, for $i\neq j$, $\sum_{i,j}(\bm{u}_{i})^{\top}\bm{u}_{j}=(\sum_{i}\bm{u}_{i})^{\top}\sum_{j}\bm{u}_{j}=0$. Using these two relations, we obtain:,
@@
Therefore, the minimum value of the incoherence parameter $\nu$ is $1+\frac{2}{n-1}$. To find the maximum value of the incoherence, using the parallelogram inequality, $\left(\bm{u}_{i}-\bm{u}_{j}\right)^{\top}\left(\bm{u}_{i}-\bm{u}_{j}\right)\leqslant 2||\bm{u}_{i}||^{2}+2||\bm{u}_{j}||^{2}\leqslant 4$. Therefore, the upper bound for $\bm{v}$ is $2\frac{n}{r}$. ∎
@@
[…]
@@
Notice that if $\|\bm{D}\|_{\infty}=\mathcal{O}(r^{2})$ and $\lambda_{r}(\bm{X})=\mathcal{O}(n)$, then $\nu=\mathcal{O}(r)$. As $r$ is most frequently either 222 or 333, this implies $\nu=\mathcal{O}(1)$ for relevant datasets, which is assumed throughout this work. We will now show that data drawn from bounded isotropic distributions exhibits this property.
@@
[[49, Page 31]] Let $\{\bm{p}_{i}\}_{i=1}^{n}\sim\mu$ where $\mu$ is a probability measure defined on $\mathbb{R}^{r}$, and let $\bm{P}=[\bm{p}_{1}\,\cdots\,\bm{p}_{n}]^{\top}\in\mathbb{R}^{n\times r}$. Define the covariance matrix of $\mu$ as $\bm{\Sigma}$. If $n\geq C(t/\varepsilon)^{2}r$ for some constant $C>0$, then with probability at least $1-2\exp(-t^{2}n)$
@@
This result is a simple application of the Hanson-Wright inequality, seen in Theorem A.3. First, for column vectors $\bm{p}_{i},\bm{p}_{j}$, notice that
@@
Previously, we have shown that $\mathbb{E}\left[\left(\bm{p}_{i}-\bm{p}_{j}\right)^{\top}\left(\bm{p}_{i}-\bm{p}_{j}\right)\right]=2r$. Furthermore, $\|\bm{A}\|_{\mathrm{F}}^{2}=4r$, and $\|\bm{A}\|\leq 2$ by Gershgorin’s circle theorem. The result follows from an application of Theorem A.3. ∎
@@
We can now see that $\|\bm{D}\|_{\infty}\leq 2r+4C^{2}\sqrt{r}\log{n}$ with high probability for a sub-Gaussian isotropic distribution. Furthermore, from Lemma 3.2 we know that $\lambda_{r}(\bm{X})=cn$ for some $\mathcal{O}(1)$ constant $c>0$. As such, we can see that the incoherence constant can be upper-bounded using (16) by
@@
We note that the analysis in this section exclusively pertained to data generated from isotropic measures. These techniques can be extended to centered and bounded anisotropic sub-Gaussian measures, and one can show the resulting bound for $\nu=\mathcal{O}(\kappa\log{n}/\sqrt{r})$, where $\kappa$ is the condition number of $\bm{X}$. We provide a proof of this result in Lemma F.3.
@@
As such, standard incoherence in the EDG framework represents a re-scaled $\ell_{2}$ norm maximum of the underlying point cloud.
3.1 Finer Interpretation of EDG Incoherence and Applications
Throughout this work, we have treated incoherence as an index-by-index bound; that is to say that we only consider terms such as $\|\mathcal{P}_{U}\bm{w}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}$. We wish to investigate this in more detail now. The main technical problem that the incoherence assumption provides a solution for is in the variance estimations used in concentration inequalities, such as in Theorem 5.3, for example. This variance estimate comes from a Gershgorin style upper bound on the matrix $\tilde{\bm{H}}=[\langle\mathcal{P}_{U}\bm{w}_{\bm{\alpha}},\mathcal{P}_{U}\bm{w}_{\bm{\beta}}\rangle]\in\mathbb{R}^{L\times L}$, seen in Lemma A.6. The eigenvalue bound leverages the fact that, if $\bm{\alpha}\cap\bm{\beta}=\emptyset$, $\langle\mathcal{P}_{U}\bm{w}_{\bm{\alpha}},\mathcal{P}_{U}\bm{w}_{\bm{\beta}}\rangle=0$, and the other terms we use Assumption 5.1 in tandem with Cauchy-Schwarz to get a uniform bound on the non-zero entries. This yields an upper bound that is used to estimate the variance term in the concentration inequalities. We argue here that a more fine-grained representation of incoherence could potentially sharpen incoherence results and lead to more geometrically-optimal sampling strategies in the future.
For the Gershgorin estimate, we need to estimate $|\langle\mathcal{P}_{U}\bm{w}_{\bm{\alpha}},\mathcal{P}_{U}\bm{w}_{\bm{\beta}}\rangle|$ for all non-zero entries of $\tilde{\bm{H}}$. Without loss of generality, we assume that $\bm{\alpha}=(i,j)$ and $\bm{\beta}=(i,k)$ for $i,j,k\in[n]$. Following a nearly identical chain of computations as in Remark 4, one can show that
@@ 3.1 Finer Interpretation of EDG Incoherence and Applications
This interpretation indicates that what might be more relevant to variance minimization is sampling more orthogonal angles with respect to a whitened dataset, rather than just considering lengths. This could lead to more optimal non-uniform sampling techniques for solving the EDG problem.
4 The Riemannian Dual Basis Approach to EDG
With the goal of translating the standard matrix completion problem to Gram matrix completion of a ground truth matrix $\bm{X}\in\mathbb{S}$, where $\mathbb{S}=\{\bm{Y}\in\mathbb{R}^{n\times n}|\bm{Y}=\bm{Y}^{T},\bm{Y}\bm{1}=\bm{0}\}$, the most direct adaptation of the work conducted in [51] would be defining an objective function by analogy to (27) as follows:
@@ 4 The Riemannian Dual Basis Approach to EDG
where $C_{\bm{\alpha}\bm{\alpha}}=p$ for all $\bm{\alpha}$, and $C_{\bm{\alpha}\bm{\beta}}=1$ for all $\bm{\alpha}\neq\bm{\beta}$. This diagonal re-scaling is introduced to make sure that $\mathbb{E}[\mathcal{M}_{\Omega}]=p^{2}\mathcal{I}$. Previous literature introduced an unscaled form of this operator, i.e. $C_{\bm{\alpha}\bm{\alpha}}=1$, computed as $\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}$ [48]. This operator does not concentrate around the identity operator, demonstrated in Lemma B.1, and as such a re-scaled form of the operator must be considered. The new operator is self-adjoint, and as such we can define the following objective function for the EDG problem using this operator:
@@
Algorithm 1 Dual Basis approach to Distance Geometry using Riemannian Gradient Descent (DB-DG-RGD)
1:Initialization: $\bm{X}_{0}=\bm{U}_{0}\bm{D}_{0}\bm{U}_{0}^{\top}$
2:for $l=0,1,\cdots$ do
3:  $\bm{G}_{l}=\mathcal{M}_{\Omega}(\bm{X}-\bm{X}_{l})$
4:  $\alpha_{l}=\frac{\left\|\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}\right\|_{\mathrm{F}}^{2}}{\langle\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l},\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}\rangle}$
5:  $\bm{W}_{l}=\bm{X}_{l}+\alpha_{l}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}$
6:  $\bm{X}_{l+1}=\mathcal{H}_{r}(\bm{W}_{l})$
7:end for
8:Output: $\bm{X}_{\mathrm{rec}}$, Gram matrix estimate after $k$ iterations
In the approach seen in Algorithm 1, the thin spectral decomposition in the gradient descent scheme is the most expensive, especially when $n$ is large. As described previously, the authors in [51] found an efficient way to reduce the computational complexity of this decomposition from $\mathcal{O}(rn^{2})$ to $\mathcal{O}(r^{3})+\mathcal{O}(nr^{2})$, substantially reducing the cost per iteration, which we implement as well. We note that in Algorithm 1 the reconstruction of the ground truth Gram matrix $\bm{X}$ is equivalent to the reconstruction of $\bm{D}$, as there is a one-to-one correspondence between $\bm{X}$ and $\bm{D}$ through (2).
@@
$\mathcal{S}_{\Omega}^{\ast}\bm{H}^{-1}\mathcal{S}_{\Omega}=\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}$
As $\mathcal{F}_{\Omega}$ exhibits the desired concentration properties (see Lemma B.6) but does not become the identity when $\Omega=\mathbb{I}$, this motivated the investigation into $\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}$. Further investigation in Lemma B.1 validates the necessity of considering a rescaled variant of $\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}$ to ensure concentration around $\mathcal{I}$, resulting in $\mathcal{M}_{\Omega}$. In essence, the terms associated with $\bm{\alpha}\neq\bm{\beta}$ are symmetrizing terms, and the rescaling $p$ for the $\bm{\alpha}=\bm{\beta}$ terms are debiasing terms.

### v2

@@ 2.2 Dual Basis Approach to EDMC
It can also easily be verified that the dimension of the linear space $\mathbb{S}$ is L=n(n1)/2L=n(n-1)/2. Using this basis, we can realize each entry of the squared distance matrix as the trace inner product of the Gram matrix with the basis. Formally, $D_{ij}=\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle$ for $\bm{\alpha}=(i,j)$. Further, we can introduce the dual basis to $\{\bm{w}_{\bm{\alpha}}\}$, denoted as $\{\bm{v}_{\bm{\alpha}}\}$, and represent any centered Gram matrix $\bm{X}$ using the following expansion:
@@
The advantage of the dual basis representation is that it allows us to recast the EDMC problem as a low-rank matrix recovery problem where we observe a subset of the expansion coefficients. In [28], this dual basis formulation has been used to provide theoretical guarantees for the convex program given in (5).
@@
From the bi-orthogonality relationship of the dual basis, it follows that $\mathcal{R}_{\Omega}^{2}=\mathcal{R}_{\Omega}$ if $\Omega$ does not have repeated indices. The adjoint of $\mathcal{R}_{\Omega}$ admits the following form:
@@
3 Geometric Interpretation of EDMC Incoherence
In pathological cases, the ground truth matrix $\bm{X}$ may exhibit a sparse representation in the basis $\{\bm{w}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$, which could lead to challenges in its recovery from sampled measurements. While the concept of incoherence is well-established in the standard matrix completion literature, the condition specific to the EDMC problem slightly differs in structure and admits a natural geometric interpretation. This section is devoted to a detailed examination of this geometric perspective. We will state more formally the incoherence assumptions in Section 5, but we will first introduce one of the conditions below. We say that a rank-$r$ Gram matrix $\bm{X}\in\mathbb{R}^{n\times n}$ is $\nu$-incoherent with respect to $\{\bm{w}_{\bm{\alpha}}\}_{\bm{\alpha}\in\mathbb{I}}$ if the following statement holds:
@@ 3 Geometric Interpretation of EDMC Incoherence
We remark that the above is inspired by the standard incoherence condition, which up to a scaling factor, states that
@@
The incoherence condition introduced in (11) can be interpreted in terms of the underlying point cloud data. For the specific case of the EDMC problem, (11) can be expanded as follows for $\bm{\alpha}=(i,j)$ with $i<j$:
@@
We consider $\sum_{(i,j):i<j}\left(\bm{u}_{i}-\bm{u}_{j}\right)^{\top}\left(\bm{u}_{i}-\bm{u}_{j}\right)$. Note that $\sum_{i}(\bm{u}_{i})^{\top}\bm{u}_{i}=\mathrm{Trace}(\bm{U}\bm{U}^{\top})=\mathrm{Trace}(\bm{U}^{\top}\bm{U})=r$. Since we assume centered configurations, $\bm{U}^{\top}\bm{1}=\bm{0}$. It then follows that, for $i\neq j$, $\sum_{i,j}(\bm{u}_{i})^{\top}\bm{u}_{j}=(\sum_{i}\bm{u}_{i})^{\top}\sum_{j}\bm{u}_{j}=0$. Using these two relations, we obtain:
@@
The above equality notes that the sum of $L$ terms is $(n+1)r$. Therefore, the maximum summand must be at least $\frac{(n+1)r}{L}$. In particular, we have:
@@
We first start our interpretation for the case where $\bm{\Lambda}$ is the identity matrix. In this setting, for any pair $(i,j)$, the expression $\left(\bm{p}_{i}-\bm{p}_{j}\right)^{\top}\left(\bm{p}_{i}-\bm{p}_{j}\right)$ is the squared Euclidean distance between the points $\bm{p}_{i}$ and $\bm{p}_{j}$. Hence, the incoherence can be directly linked to the maximum distance among the points. We now provide an interpretation of (14) in the general case. The quantity therein suggests that incoherence serves as a measure of how the displacement vectors $\bm{p}_{i}-\bm{p}_{j}$ align with the principal components of the embedding. In particular, for a fixed choice of $\bm{\Lambda}$, varying the matrix $\bm{U}$ leads to different sets of points. If the displacement vectors tend to align with directions corresponding to the smallest principal components (i.e., those with the lowest variance), the incoherence is expected to be high. Conversely, if they align more with the dominant components (those with the highest variance), the incoherence tends to be low. In essence, high incoherence indicates that certain pairs of points are stretching significantly in directions where the embedding space has low variance.
@@
Notice that if $\|\bm{D}\|_{\infty}=\mathcal{O}(r^{2})$ and $\lambda_{r}(\bm{X})=\mathcal{O}(n)$, then $\nu=\mathcal{O}(r)$. As $r$ is most frequently either $2$ or $3$, this implies $\nu=\mathcal{O}(1)$ for relevant datasets, which is assumed throughout this work. We will now show that data drawn from bounded isotropic distributions exhibits this property.
@@
[[52, Page 31]] Let $\{\bm{p}_{i}\}_{i=1}^{n}\sim\mu$ where $\mu$ is a probability measure defined on $\mathbb{R}^{r}$, and let $\bm{P}=[\bm{p}_{1}\,\cdots\,\bm{p}_{n}]^{\top}\in\mathbb{R}^{n\times r}$. Define the covariance matrix of $\mu$ as $\bm{\Sigma}$. If $n\geq C(t/\varepsilon)^{2}r$ for some constant $C>0$, then with probability at least $1-2\exp(-t^{2}n)$
@@
This result is a simple application of the Hanson-Wright inequality, seen in A.3. First, given vectors $\bm{p}_{i},\bm{p}_{j}$, notice that
@@
Previously, we have shown that $\mathbb{E}[\left(\bm{p}_{i}-\bm{p}_{j})^{\top}(\bm{p}_{i}-\bm{p}_{j})\right]=2r$. Furthermore, $\|\bm{A}\|_{\mathrm{F}}^{2}=4r$, and $\|\bm{A}\|\leq 2$ by Gershgorin circle theorem. The result follows from an application of A.3. ∎
@@
We can now see that $\|\bm{D}\|_{\infty}\leq 2r+4C^{2}\sqrt{r}\log{n}$ with high probability for a sub-Gaussian isotropic distribution. Furthermore, from 3.2, we know that $\lambda_{r}(\bm{X})=cn$ for some $\mathcal{O}(1)$ constant $c>0$. As such, we can see that the incoherence constant can be upper-bounded using (16) by
@@
We note that the analysis in this section exclusively pertained to data generated from isotropic measures. These techniques can be extended to centered and bounded anisotropic sub-Gaussian measures, and one can show the resulting bound for $\nu=\mathcal{O}(\kappa\log{n}/\sqrt{r})$, where $\kappa$ is the condition number of $\bm{X}$. We provide a proof of this result in F.3.
@@
As such, standard incoherence in the EDMC framework represents a re-scaled $\ell_{2}$ norm maximum of the underlying point cloud.
3.1 Finer Interpretation of EDMC Incoherence and Applications
Throughout this work, we have treated incoherence as an index-by-index bound; that is to say that we only consider terms such as $\|\mathcal{P}_{U}\bm{w}_{\bm{\alpha}}\|_{\mathrm{F}}^{2}$. We wish to investigate this in more detail now. The main technical problem that the incoherence assumption provides a solution for is in the variance estimations used in concentration inequalities, such as in 5.3, for example. This variance estimate comes from a Gershgorin style upper bound on the matrix $\tilde{\bm{H}}=[\langle\mathcal{P}_{U}\bm{w}_{\bm{\alpha}},\mathcal{P}_{U}\bm{w}_{\bm{\beta}}\rangle]\in\mathbb{R}^{L\times L}$, seen in A.6. The eigenvalue bound leverages the fact that, if $\bm{\alpha}\cap\bm{\beta}=\emptyset$, $\langle\mathcal{P}_{U}\bm{w}_{\bm{\alpha}},\mathcal{P}_{U}\bm{w}_{\bm{\beta}}\rangle=0$, and the other terms we use 5.1 in tandem with Cauchy-Schwarz to get a uniform bound on the non-zero entries. This yields an upper bound that is used to estimate the variance term in the concentration inequalities. We argue here that a more fine-grained representation of incoherence could potentially sharpen incoherence results and lead to more geometrically-optimal sampling strategies in the future.
For the Gershgorin estimate, we need to estimate $|\langle\mathcal{P}_{U}\bm{w}_{\bm{\alpha}},\mathcal{P}_{U}\bm{w}_{\bm{\beta}}\rangle|$ for all non-zero entries of $\tilde{\bm{H}}$. Without loss of generality, we assume that $\bm{\alpha}=(i,j)$ and $\bm{\beta}=(i,k)$ for $i,j,k\in[n]$. Following a nearly identical chain of computations as in Remark Remark 4, one can show that
@@ 3.1 Finer Interpretation of EDMC Incoherence and Applications
This interpretation indicates that what might be more relevant to variance minimization is sampling more orthogonal angles with respect to a whitened dataset, rather than just considering lengths. This could lead to more optimal non-uniform sampling techniques for solving the EDMC problem.
4 The Riemannian Dual Basis Approach to EDMC
With the goal of translating the standard matrix completion problem to Gram matrix completion of a ground truth matrix $\bm{X}\in\mathbb{S}$, where $\mathbb{S}=\{\bm{Y}\in\mathbb{R}^{n\times n}|\bm{Y}=\bm{Y}^{\top},\bm{Y}\bm{1}=\bm{0}\}$, the most direct adaptation of the work conducted in [54] would be defining an objective function by analogy to (28) as follows:
@@ 4 The Riemannian Dual Basis Approach to EDMC
where $C_{\bm{\alpha}\bm{\alpha}}=p$ for all $\bm{\alpha}$, and $C_{\bm{\alpha}\bm{\beta}}=1$ for all $\bm{\alpha}\neq\bm{\beta}$. This diagonal re-scaling is introduced to make sure that $\mathbb{E}[\mathcal{M}_{\Omega}]=p^{2}\mathcal{I}$. Previous literature introduced an unscaled form of this operator, i.e. $C_{\bm{\alpha}\bm{\alpha}}=1$, computed as $\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}$ [51]. This operator does not concentrate around the identity operator, demonstrated in B.1, and as such a re-scaled form of the operator must be considered. The new operator is self-adjoint, and as such we can define the following objective function for the EDMC problem using this operator:
@@
Algorithm 1 Dual Basis Riemannian EDMC (DBRE)
1:Input: $\Omega$, $\{\langle\bm{X},\bm{w}_{\bm{\alpha}}\rangle\}_{\bm{\alpha}\in\Omega}$, $\mathrm{rank}(\bm{X})=r$
2:Initialization: $\bm{X}_{0}=\bm{U}_{0}\bm{D}_{0}\bm{U}_{0}^{\top}$
3:for $l=0,1,\cdots$ do
4:  $\bm{G}_{l}=\mathcal{M}_{\Omega}(\bm{X}-\bm{X}_{l})$
5:  $\alpha_{l}=\frac{\left\|\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}\right\|_{\mathrm{F}}^{2}}{\langle\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l},\mathcal{M}_{\Omega}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}\rangle}$
6:  $\bm{W}_{l}=\bm{X}_{l}+\alpha_{l}\mathcal{P}_{\mathbb{T}_{l}}\bm{G}_{l}$
7:  $\bm{X}_{l+1}=\mathcal{H}_{r}(\bm{W}_{l})$
8:end for
9:Output: $\bm{X}_{\mathrm{rec}}$: Estimated Gram matrix.
In the approach seen in Algorithm 1, the thin spectral decomposition in the gradient descent scheme is the most expensive, especially when $n$ is large. As described previously, the authors in [54] found an efficient way to reduce the computational complexity of this decomposition from $\mathcal{O}(rn^{2})$ to $\mathcal{O}(r^{3}+r^{2}n)$, substantially reducing the cost per iteration, which we implement as well. We note that in Algorithm 1 the reconstruction of the ground truth Gram matrix $\bm{X}$ is equivalent to the reconstruction of $\bm{D}$, as there is a one-to-one correspondence between $\bm{X}$ and $\bm{D}$ through (2).
@@
$\mathcal{S}_{\Omega}^{\ast}\bm{H}^{-1}\mathcal{S}_{\Omega}=\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}.$
As $\mathcal{F}_{\Omega}$ exhibits the desired concentration properties (see B.6) but does not become the identity when $\Omega=\mathbb{I}$, this motivated the investigation into $\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}$. Further investigation in B.1 validates the necessity of considering a rescaled variant of $\mathcal{R}^{\ast}_{\Omega}\mathcal{R}_{\Omega}$ to ensure concentration around $\mathcal{I}$, resulting in $\mathcal{M}_{\Omega}$.